OpenStudy (anonymous):

For the following equation, state the value of the discriminant and then describe the nature of the solutions. 10x^2+10x-19=0 What is the value of the discriminant ? Which one of the statements below is correct? A) The equation has two real solutions. B) The equation has one real solution. C)The equation has two imaginary solutions.

OpenStudy (anonymous):

\[b^2-4ac=10^2-4\times 10\times (-19)\]which is some positive number i am too lazy to compute. but because it is positive that means choice A